1. The square of an interger is 30 more than the integer. Find the integer.
I tried.... x^2 = x+30 to get a solution of x=6 and x-5
I am unsure if this is the right equation for the problem above.
2. The sum of an integer and its square is 30. Find the number.
I tried..... x + x^2= 30 to get a solution of x=-6 and x=5
Once again I am unsure if this is the right equation for the problem above.

Answer by Edwin McCravy(20054)

1. The square of an interger is 30 more than the integer. Find the integer.
I tried.... x^2 = x+30 to get a solution of x=6 and x-5
I am unsure if this is the right equation for the problem above.
 

Then by all means learn to check your solutions.  Remember, when you
check a WORD problem, you check each solution in the WORDS, never in 
the equation, for you might have the right answer to the wrong 
equation!

Here, let's go a-checking. It's easy.

The first sentence says:

>>...The square of an integer is 30 more than the integer...<<

Check that with 6. In other words, is it true that:

the square of 6 is 30 more than 6?

the square of 6 is 36, and lo and behold, 36 really is 30 
more than 6. !!!

So we know that 6 is correct.

But we still have another solution to check, x=-5

Again, the first sentence says:

>>...The square of an integer is 30 more than the integer...<<

Check that now with -5. In other words, is it true that:

the square of -5 is 30 more than 6.

the square of -5, or (-5)² is +25 and, holy smokes, +25 
really is 30 more than -5. !!!

So we know that -5 is also correct.

---------------------------------------------


2. The sum of an integer and its square is 30. Find the number.
I tried..... x + x^2= 30 to get a solution of x=-6 and x=5
 
Then you should check these two solutions.

The first sentence says:

>>...The sum of an integer and its square is 30...<<

Check that with -6. In other words, is it true that:

the sum of -6 and its square is 30?

The square of -6 is +36 and heyyy, the sum of -6 and 36 
really is 30. !!!

So we know that -6 is correct.

But we still have another solution to check, x=5

So we check that with 5. In other words, is it true that:

the sum of 5 and its square is 30?

The square of 5 is 25 and whaddayaknow, the sum 
of 5 and 25 really is 30. !!!

So we know that 5 is correct.

See how easy it is to check a number back into the 
words to see if it is the correct answer?

---------------------------------------------
Edwin